OpenStudy (anonymous):

Petrolyn motor oil is a combination of natural oil and synthetic oil. It contains 3 liters of natural oil for every 2 liters of synthetic oil. If 198 liters of synthetic oil are used, how many liters of Petrolyn oil will be made?

OpenStudy (anonymous):

I bet you are annoyed at this point of me but I just have this big packet of homework and I need help. @jim_thompson5910 If you can help thanks and if you cant its fine I understand

jimthompson5910 (jim_thompson5910):

so we'll have a proportion we need to set up first do you have a proportion? or are you stuck on that part?

OpenStudy (anonymous):

Im stuck on the proportion

OpenStudy (anonymous):

3/2=x/198?

jimthompson5910 (jim_thompson5910):

this will be the basic template \[\Large \frac{\text{amount of natural oil}}{\text{amount of synthetic oil}} = \frac{\text{amount of natural oil}}{\text{amount of synthetic oil}}\]

OpenStudy (anonymous):

so then it would be 2x=594?

jimthompson5910 (jim_thompson5910):

we know that " It contains 3 liters of natural oil for every 2 liters of synthetic oil" focus on one side, say the right side replace "natural oil" with 3 replace "synthetic oil" with 2 \[\Large \frac{\text{amount of natural oil}}{\text{amount of synthetic oil}} = \frac{\text{amount of natural oil}}{\text{amount of synthetic oil}}\] \[\Large \frac{3}{2} = \frac{\text{amount of natural oil}}{\text{amount of synthetic oil}}\]

jimthompson5910 (jim_thompson5910):

yeah you have the right idea, doing 2 more replacements gives \[\Large \frac{3}{2} = \frac{\text{amount of natural oil}}{\text{amount of synthetic oil}}\] \[\Large \frac{3}{2} = \frac{x}{198}\]

jimthompson5910 (jim_thompson5910):

and yes, when you cross multiply you will get 2x = 594

OpenStudy (anonymous):

right yes so then x would be

OpenStudy (anonymous):

x=297

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

Thank you again

jimthompson5910 (jim_thompson5910):

297 isn't the amount of Petrolyn oil though 297 is the amount of natural oil

OpenStudy (anonymous):

so we addor?

jimthompson5910 (jim_thompson5910):

hmm let me think

jimthompson5910 (jim_thompson5910):

oh I'm overthinking, yeah we add amount of petrolyn oil = (amount of natural oil) + (amount of synthetic oil)

OpenStudy (anonymous):

so its 495

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

great! :)